The area enclosed by the curve y2 + x4 = x2 is-
(a)2/3
(b)4/3
(c)8/3
(d)10/3
ย 

Grade:12

1 Answers

Arun
5 years ago
Dear student
ย 
This curve is symmetrical along the x and y axis, so we can integrate the area in the first quadrant and multiply it by 4.ย 

y = x(โˆš(1-x^2))ย 

A = โˆซ 0->1 x(โˆš(1-x^2)) dxย 

A = -1/3 (1-x^2)^(3/2) | 0->1ย 

A = 1/3ย 

Total Area is 1/3 x 4 = 4/3